### v2

@@ 3.2 The benchmark
Dataset Nodes Links $Q(\text{parties})$ $Q(\text{coalitions})$
@@
To compare different algorithms not just based on the small minority of users that we manually annotated, we extended the annotations to the rest of the network using a well-known Label propagation algorithm [42] –taking the most frequent label out of 100 independent realizations. These propagated labels cannot be considered equally accurate, but they arguably represent the best possible guess of a user’s political leaning, which can be made only based on how that user is connected to verified politicians in the retweet network.
Let us notice that such an approach favors our models, in principle, since we are using the same label propagation algorithm to label standard users. In a sense, if we correctly tag $\top$ users, we will implicitly also meet the benchmark for $\bot$ users. It is indeed true, but it is an honest way to propagate the labels of verified accounts, as we cannot trust, in general, the one of other accounts that could be present in the user description.
3.3 Limits of standard methods
Let us start by considering the partitions obtained running four off-the-shelf community-detection algorithms on the entire retweet network: (i) Louvain, a heuristic for modularity maximization; (ii) Label propagation, a message passing algorithm where each node tries to align with the majority of its neighbors; (iii) Infomap, which is based on finding the optimal encoding for random walks on the graph; (iv) Inference, whose rationale is inferring the parameters of a stochastic block-model through a maximum-likelihood approach.
For Louvain, Label propagation and Infomap, we used the implementation available in igraph (igraph.org, Python interface). For Inference, we used the minimize_blockmodel_dl() function provided in graph-tool (graph-tool.skewed.de), that fits a stochastic block-model to the given graph by minimizing the model’s description length. To guide the inference process, we set it explicitly to search for a planted-partition, that is, exclusively for assortative communities. In addition, we refined the algorithm as recommended in graph-tool’s documentation, that is, by running 100 times the multiflip_mcmc_sweep() method with zero temperature and 100 iterations.
We ran each algorithm 100 times [^4] Each single run of the Louvain algorithm is already obtained taking the partition with the greatest modularity out of 100 independent executions with shuffled vertices. Similarly, each run of the Label propagation algorithm is already obtained by taking, for each node, the most frequent label out of 100 independent executions. and measured: the $\mathrm{VM}_{1}$ between each partition, restricted to the annotated users, and the annotations, both at the party and coalition level; the pairwise $\mathrm{VM}_{1}$ between partitions obtained with different algorithms, restricted to the verified users. Fig. 3 reports these $\mathrm{VM}_{1}$ values, considering only the parties annotations. We focused on party annotations, but similar results are observed with other algorithms and with coalitions.
Fig. 3 highlights the limited accuracy in recognizing the affiliation of renowned politicians –$\mathrm{VM}_{1}\in[0.4,0.6]$– of the communities assigned to verified users by different off-the-shelf algorithms.
Furthermore, as shown in Fig. 10 in Appendix C, standard algorithms show great variability of the communities assigned to verified users by different off-the-shelf algorithms –pairwise $\mathrm{VM}_{1}\in[0.3,0.7]$. Overall, these findings motivate the research for a different approach to the detection of discursive communities on Twitter/X.
@@ 3.3 Limits of standard methods
Figure 3: Limits of standard methods. The $\mathrm{VM}_{1}$ between partitions obtained with off-the-shelf algorithms, restricted to the annotated users, and the parties’ annotations. Standard methods show a limited accuracy in recognising the affiliation of renowned politicians.
3.4 $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$
We now show the results of our methods, i.e. $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, when we choose verified users as the class of content creators: in this case, we will write $\top=V$. In the following, the communities on the network of $V$ users are obtained by running the Louvain algorithm on the projected network and are then propagated to the rest of the network. We tested our community detection algorithm, as the Label propagation and Infomap: their results are very similar to those obtained with Louvain ($\mathrm{VM}_{1}>0.8$) and yield similar results also for both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$. Communities obtained with the inference-based method are instead way less accurate in reconstructing the annotations.
3.4.1 Comparison with standard methods
In Fig. 4, we use the $\mathrm{VM}_{1}$ to show how accurately each method reconstructs the manual annotations at the party and coalition level, respectively in the first and second row. Each violin represents the distribution over 100 independent runs, and we compare $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with the partitions obtained running the Louvain algorithm on the entire retweet network –which performs no worse than other algorithms (cfr. Fig. 3). $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ systematically outperform Louvain, with $\mathrm{MonoDC}_{V}$ generally preferable in identifying parties while $\mathrm{BiDC}_{V}$ in identifying coalitions. The fact that in all three cases the community detection algorithm used is Louvain signifies that what ultimately matters is focusing on the right network.
For completeness, the comparison between the performances of our methods and Louvain is presented in Fig. 11 in Appendix C. Even in this case, both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ outperform Louvain. Nonetheless, let us note that, when comparing the results of the propagated labels of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}$ with the annotations of verified users propagated using the same label propagation algorithm, our methods have an advantage. Nevertheless, the performances of standard methods are worse, even when limited to annotated users.
@@ 3.4.1 Comparison with standard methods
Figure 4: Comparison of the performances of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with standard methods. The $\mathrm{VM}_{1}$ between partitions of the annotated users, obtained with different methods, and the annotations at party (first row) and coalition (second row) level. $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ systematically outperform Louvain, with $\mathrm{MonoDC}_{V}$ generally preferable in identifying parties while $\mathrm{BiDC}_{V}$ in identifying coalitions.
It could be argued that standard algorithms, when executed on the entire retweet network, return sub- or super-groups of the political parties that $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ cannot detect. Fig. 5 shows that this is not the case: when compared with manually annotated parties (top) and coalitions (bottom), $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ exhibit a consistently greater average $\mathrm{VM}_{\beta}$ than standard methods even for $\beta\neq 1$, i.e. at essentially all scales at which the political spectrum can be observed.
@@
In Fig. 6, the Sankey diagrams corresponding to each evaluated method provide a detailed depiction of the political composition of the extracted communities. $\mathrm{MonoDC}_{V}$ exhibits remarkable precision in delineating party structures —successfully distinguishing individual parties such as Azione and Italia Viva, which presented a unified list during the elections— and further reveals latent political currents within party boundaries. Conversely, $\mathrm{BiDC}_{V}$ proves highly effective in reconstructing the official electoral coalitions, achieving a near-exact correspondence with their institutional configurations. The Louvain algorithm, applied to the entire retweet network, yields a substantially noisier community structure that conflates partisan and coalition-level affiliations. Notably, Forza Italia emerges as an isolated community, distinct from its allied parties, while several political figures are clustered within communities dominated by their ideological adversaries.
(a) Louvain on RT network.
@@
Finally, since $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ capture the political leaning of users at different scales, in Fig. 7 we use the $\mathrm{VM}_{\beta}$ to check whether the clusters found by $\mathrm{MonoDC}_{V}$ are subclusters of those found by $\mathrm{BiDC}_{V}$. This seems to be true, at least in part, and especially for the two datasets, Crisis and Elections, which refer to a political phase in which the coalitions were more clearly defined.
3.4.2 Final remarks
Both $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ benefit from a pre-processing stage capable of extracting, from the original retweet data, a more informative network of just such leading content creators. The role that statistical validation plays in the two cases is, however, incomparable.
In $\mathrm{MonoDC}_{V}$, the validated retweet networks of verified accounts of politicians include 94-97% –depending on the dataset– of all edges present before the validation. The validation is performed with respect to the patterns emerged from the analysis of the entire network and the fact that almost all retweets made by politicians are statistically significant means that their activity on X is extremely focused on just the content produced by a limited niche of users. Since these users are their party’s leaders and fellow members, $\mathrm{MonoDC}_{V}$ is the approach that best captures the loyalty to the party and, hence, the difference between parties that apparently occupy a similar position in the political spectrum. The price to pay for this finer classification is its minor stability, as can be seen in Fig. 4. In this sense, $\mathrm{MonoDC}_{V}$ seems particularly promising to examine the mesoscale structure of such a subgraph to infer the main poles of the political discussion, since the phenomenon we are leveraging on is particularly strong.
Differently from the $\mathrm{MonoDC}_{V}$’s case, for $\mathrm{BiDC}_{V}$ the impact of the statistical validation is critical: the number of links in the validated network is 7-12% –depending on the dataset– of those present in the non-validated projection. When the political sides are clear, this validated signal coming from the audience returns a very stable partition (cfr. Fig. 4). In addition, $\mathrm{BiDC}_{\top}$ has the advantage that the similarity between any two content creators $u,v\in\top$ can be established even when $u$ and $v$ did not directly interact, making $\mathrm{BiDC}_{\top}$ more flexible, robust and widely usable than $\mathrm{MonoDC}_{\top}$.
For the sake of completeness, we ran the same set of community-detection algorithms considered in Section 3.3 on the non-validated network, as well as on the validated projections with different significance levels [^5] To properly target the significance level in validating the projected network, we implemented the False Discovery Rate (FDR) procedure [46], see Appendices A.1.1 and A.2.1., cfr. Fig. 12 in Appendix C. The results on the non-validated network are very poor, whereas the significance level does not seem to have a great impact on the inferred similarity network.
3.5 Other classes of content creators
In the previous subsections, we examined the behaviours of the $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ after identifying verified users as relevant content creators, i.e. the main leaders of the shared discourses. Verified users ($V$) are a very natural choice for the $\top$ layer of both $\mathrm{MonoDC}_{\top}$ –because their behavior on the platform is bound to their role outside of it– and $\mathrm{BiDC}_{\top}$ –because they are generally known to and easily categorizable by the audience.
There are other ways, however, to select users based on their actual or presumed influence on online discussions. For instance, [4] defined “influential” users those who have more followers than friends (i.e. the accounts a user follows, sometimes referred to as “followees” in the literature), and who, in the debate under study, received more mentions than those they made. Another way to measure the impact of a user in a community, commonly employed for scientists and academics, is the Hirsch index [47] –shortly, H-index– which can be adapted to X’s mode of operation by saying that a user $u$’s H-index in the debate is the largest $h$ such that at least $h$ tweets authored by $u$ received at least $h$ retweets each.
In the following, we use $I$ to denote the set of influential users and $H$ for the set of top H-index users (defined as users having H-index3absent3$\geq$ 3≥ 3, see Fig. 13 in Appendix C for a justification of the threshold). We aim to assess both what happens when $\top=I$ or $\top=H$ is used in $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, and how stable and accurate is the classification obtained for these two sets of central users when compared with the entire dataset.
@@ 3.5 Other classes of content creators
In Fig. 8, the $\mathrm{VM}_{1}$ is used to compare partitions obtained with different choices of $\top$ with the (propagated) annotations ($x$-axis). Here we consider the Elections dataset, but similar results hold for the other datasets.
In Appendix C), we also compare the cases $\top=H$ and $\top=I$ with the case $\top=V$, taken as a reference.
The choice of the $\top$ set has an impact on both $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$. Especially for $\mathrm{BiDC}_{\top}$, $\top=V$ yields the best overlap with the annotations. It could be argued that it is due to $V$ being a small set that includes all manually annotated users, but $H$ outperforms $I$ despite the former includes just 20-50% (depending on the dataset) of the annotated public figures, compared with 80-90% of the latter.
This supports the claim that $V$ users are the key to understanding the sides of the public discourse.
4 Conclusions
In this paper, we focused on X/Twitter and on so-called discursive communities, i.e. the sub-groups that naturally emerge in Online Social Networks (OSNs) as users cluster around shared narratives, language, or value systems during public debates. What we observe in OSNs is a network of interactions, where edges exist for a combination of factors – e.g. the visibility, the status and the behaviour of different users on and off the platform, and the platform’s recommendation system – and where understanding what led to the creation of a community is almost as important as detecting the community itself.
In this paper, we investigate the role of verified users in shaping discursive communities in the pre-Premium accounts era. Introduced in 2009, verified accounts were originally intended to prevent impersonation of public figures or organizations [21, 22]. At first, the platform itself selected which accounts to verify, based on their relevance in public discourse. These accounts were marked by a blue checkmark next to their name. While verification helped authenticate the presence of public figures, it also introduced a hierarchy among users—one that arguably amplified the influence of verified accounts in online debates [48].
Despite these concerns, in November 2022, following Elon Musk’s acquisition of Twitter, verification became a paid feature available through X’s Premium subscription [24].
We conduct our analysis around three key political events in Italy during 2022: the election of the President of the Republic, the political crisis, and the subsequent national elections. Our initial observation is that a strong signal of political alignment is detectable in the data: when clustering verified politicians in the retweet network by political party, the resulting partition shows high modularity. We then explore the limitations of existing community detection algorithms. When comparing their output to manually annotated data, both descriptive and inferential methods often produce noisy and imprecise community structures.
To address this, we evaluate the performance of two algorithms designed to identify discursive communities: $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$, both of which begin by labeling verified users, given their central role in public debate. The $\mathrm{MonoDC}_{V}$ approach is grounded in a typical political dynamic, where less prominent figures tend to retweet content from party leaders. Conversely, the $\mathrm{BiDC}_{V}$ method—originally proposed by [18]—builds an undirected similarity network between content creators by analyzing audience overlap. In both cases, we reduce the effect of random noise—an inherent feature of OSNs [28, 29]—by comparing empirical data to a maximum entropy null model, which accounts for user activity [30].
Both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ successfully reproduce the organization of political debates into discursive communities shaped by parties or alliances. Notably, using verified users as leading content creators consistently outperforms alternative selections based solely on user activity. This suggests that the status conferred by verification plays a more decisive role in structuring the online conversation than activity alone.
In summary, our work offers two take-home messages. First, in X/Twitter datasets from before November 2022, leveraging verified users allows for a clearer and more accurate identification of discursive communities. Alternative approaches relying solely on user activity result in noisier partitions, suggesting that user status is a major factor in shaping the online debate.
Second, our findings are based on comparisons conducted before account verification became a paid feature. This raises concerns about the potential consequences of monetized verification: by enabling users to purchase visibility and prestige, the platform may open the door to malicious actors seeking to manipulate discourse and distort the public agenda. A deeper comparison between “old” –i.e. pre-November 2022– and “new” –i.e. post-November 2022– verified users will be the target of further research.

### v3

@@ Results
Dataset Nodes in GCC Links $Q(\text{parties})$ $Q(\text{coalitions})$
@@
To extrapolate this benchmark to the broader network, we propagated community labels from verified users to the rest of the user population using a label propagation algorithm [Raghavan2007a]. While these propagated labels are inherently noisier than manual annotations, they provide a principled proxy for large-scale political alignment and enable a systematic comparison of community-detection strategies over the full dataset.
Standard community detection on the full retweet network shows limited political alignment
We applied four widely used community detection algorithms—Louvain, label propagation, Infomap, and stochastic block model inference—to the complete retweet networks, defined as the directed weighted graph including all users active during the event and all observed retweet interactions. While all methods identified non-trivial community structure, their agreement with the manual annotations was limited. V-measure scores for party affiliation ranged between 0.4 and 0.6 (cfr. Fig. 3), with substantial variability across algorithm runs.
Comparisons between algorithms further revealed significant inconsistency, with pairwise V-measures for the same dataset lying in the $[0.3,0.7]$ range (cfr. Fig. 12 in the Appendix). Qualitative inspection confirmed that standard methods often conflated ideologically distinct actors or fragmented coherent political blocs.
These results do not indicate a general failure of standard algorithms, but rather highlight the difficulty of recovering politically meaningful communities when they are applied directly to noisy retweet networks without distinguishing between heterogeneous user roles or filtering statistically expected interactions.
@@
Figure 3: Limits of standard methods. The $\mathrm{VM}_{1}$ between partitions obtained with off-the-shelf algorithms, restricted to the annotated users, and the parties’ annotations. Each violin represents the distribution over 100 independent runs. Standard methods show a limited accuracy in recognizing the affiliation of renowned politicians.
Leader-anchored and statistically validated networks improve community detection
Applying our two proposed methods, $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$, resulted in a substantial improvement in community-detection performance. When verified users were used as anchors within statistically validated networks, both methods produced partitions closely aligned with the annotated political affiliations.
$\mathrm{MonoDC}_{V}$ achieved particularly high accuracy at the party level, reflecting the focused retweet patterns through which politicians tend to amplify content from their own party leadership. By contrast, $\mathrm{BiDC}_{V}$ more effectively captured coalition-level structures, as it relies on shared audience behaviour rather than direct interactions among leaders.
Quantitatively, V-measure scores for both methods systematically exceeded those of standard algorithms across all datasets. $\mathrm{MonoDC}_{V}$ reached values above 0.75 for party-level annotations, while $\mathrm{BiDC}_{V}$ achieved comparably strong results at the coalition level (cfr. Fig. 4). These findings remained robust across different values of the resolution parameter $\beta$ (cfr. Fig. 5) and when considering propagated labels on the full network (cfr. Fig. 13 in the Appendix).
As a baseline, we apply the Louvain algorithm to the complete retweet network. This representation includes neither statistical filtering nor a distinction between verified and non-verified users, and thus reflects the raw interaction structure dominated by highly heterogeneous activity patterns.
Although $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ also rely on Louvain for the final partitioning step, the comparison serves a different purpose. In our framework, Louvain is applied after maximum-entropy filtering and on networks anchored to a coherent set of political leaders. Running Louvain directly on the raw retweet network therefore isolates the contribution of preprocessing and leader selection, rather than acting as a competing methodological alternative. Notably, the Louvain algorithm performs no worse than the others in recovering the annotations (cfr. Fig. 3).
Overall, these results indicate that verified accounts function as effective anchors for political discursive communities within a statistically validated framework. The improved performance does not stem from verified status alone, but from its combination with maximum-entropy filtering and with the institutional coherence of the verified-user set under the pre-2022 verification regime.
@@
Figure 4: Comparison of the performances of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with standard methods. The $\mathrm{VM}_{1}$ between partitions of the annotated users, obtained with different methods, and the annotations at party (first row) and coalition (second row) level. Each violin represents the distribution over 100 independent runs, and we use the Louvain algorithm as a benchmark because it performs no worse than other algorithms (cfr. Fig. 3). $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ systematically outperform Louvain, with $\mathrm{MonoDC}_{V}$ generally preferable in identifying parties while $\mathrm{BiDC}_{V}$ in identifying coalitions.
@@
MonoDC captures party structure, while BiDC reveals coalitional organization
The complementary strengths of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ arise from the nature of their respective network constructions. $\mathrm{MonoDC}_{V}$ relies on direct retweet interactions among verified users, which tend to encode endorsement and intra-party signalling. As a result, its statistically validated networks retain a large fraction of edges (94–97%), indicating dense and coherent communication among political actors.
$\mathrm{BiDC}_{V}$, by contrast, infers similarity among leaders through audience overlap. After statistical validation, only 7–12% of edges are retained, reflecting the fact that only a small subset of shared audiences carries meaningful political alignment. This filtering allows $\mathrm{BiDC}_{V}$ to surface stable coalition-level structures, particularly in phases where electoral alliances are clearly defined.
Consistently, the $\mathrm{VM}_{\beta}$ between the partitions produced by the two methods shows that $\mathrm{MonoDC}_{V}$ communities tend to be subclusters of those identified by $\mathrm{BiDC}_{V}$, especially for the Crisis and Elections datasets (cfr. Fig. 6).
A qualitative comparison further illustrates these differences (cfr. Fig. 7). $\mathrm{MonoDC}_{V}$ yields fine-grained partitions aligned with individual parties, whereas $\mathrm{BiDC}_{V}$ aggregates these into broader coalitional structures. In contrast, Louvain applied to the full retweet network produces a noisier mapping that conflates party- and coalition-level affiliations.
We attribute this fragmentation not to political complexity, but to the structural noise of the raw retweet network, which is substantially reduced by maximum-entropy filtering.
(a) Louvain on the retweet network.
@@
Activity-based definitions of leaders yield weaker political partitions
[…]
In all cases, performance deteriorated relative to anchoring communities on verified users (cfr. Fig. 8). While activity-based criteria identify highly active accounts, their discursive behavior is more heterogeneous and less predictive of political alignment. In particular, H-index-based selections include journalists and commentators whose audiences span ideological boundaries, blurring community structure.
These results highlight that the effectiveness of our framework depends not only on selecting active users, but on identifying content creators whose publicly recognized roles constrain and stabilize their online behavior. Further comparison can be found in Section D of the Appendix.
@@
Discussion
Digital platforms play an increasingly influential role in shaping the structure and visibility of public discourse. One crucial contribution of this work is the characterization of a general methodology for extracting meaningful discursive communities from noisy social data. We introduced a leader-based framework that explicitly accounts for the asymmetric structure of online participation. By separating content producers from their audiences, our approach allows communities to emerge more clearly than methods that treat all users symmetrically. We showed that identifying and positioning content producers is a necessary first step, after which the broader user base can be reliably assigned.
[…]
In addition, through a systematic comparison of three alternative criteria for selecting leading users, our findings demonstrate that institutional status, rather than raw activity, provides the strongest signal for organizing online debates. We found, in fact, that pre-2022 verified accounts represent the most informative anchors for community detection. We interpret this result as a consequence of their public recognizability and the behavioral constraints imposed by their off-platform roles, rather than as an intrinsic property of verification itself.
While the institutional meaning of verification has changed following the introduction of paid subscriptions, the methodological insights of this study remain valid. In particular, our results suggest that carefully curated sets of publicly relevant actors–identified either through institutional roles or external information–can serve as effective seeds for analyzing online discourse, at least in the well-studied and highly structured context of political debates. This insight opens the door to more efficient data-collection strategies, in which the ego networks of a limited number of high-relevance users are sufficient to recover much of the political structure of online debates, even under restrictive platform access conditions. Indeed, similar strategies have been implemented, for instance in [Brugnoli2024]. More broadly, our study reinforces the idea that the organization of digital public debate potentially depends not only on individual user behavior, but also on how platforms encode and signal authority. While the methods proposed here are technical, they contribute to a growing body of work showing that discursive structure, legitimacy, and visibility are deeply intertwined in online public spheres.
As any other product of research, our analyses have some limitations. First, our empirical results pertain exclusively to the pre-2022 verification regime, under which verified accounts corresponded to publicly recognized figures and formed coherent, politically aligned communities. In this historical setting, verification acted as a reliable proxy for public relevance and institutional embeddedness, providing a stable anchor for discursive structures. The transition to paid verification introduces a qualitatively different regime, whose effects fall outside the scope of the present analysis. Concerns about potential distortions of online discourse–such as increased visibility for actors lacking public accountability or the erosion of previously coherent community structures–should therefore be interpreted as speculative and grounded in plausible mechanisms rather than in direct empirical evidence. Whether political discourse under the new regime will remain similarly structured, fragment into less interpretable patterns, or reorganize around alternative signals of authority remains an open empirical question. Future research should explicitly address this gap by analyzing data collected after the introduction of paid verification, enabling direct comparisons between the two regimes. Such studies could assess whether the structural role once played by verified users has weakened, persisted, or been replaced by other indicators of public relevance.
Second, our empirical analysis focuses on political debates within a single national context, Italy, and on a single social media platform, Twitter/X. This choice reflects both data availability and methodological considerations: analyzing a political system that is well known to the authors allows for more reliable manual annotations and a more accurate interpretation of political actors, coalitions, and discursive dynamics. Importantly, this focus does not represent an intrinsic limitation of the proposed framework, which is grounded in structural features of online discourse–such as strong participation asymmetries and the presence of discussion leaders–that have been widely documented across platforms, languages, and cultural settings [Watts2007, Hilbert2017, Bracciale2018]. Consistently with this perspective, previous studies have already shown that $\mathrm{BiDC}$ performs well in different countries [Mattei2022, Declerck2022a, DeClerck2024], while the present work provides the first systematic investigation of $\mathrm{MonoDC}$ and a direct comparison between alternative leader-based constructions. Moreover, the proposed approach does not rely on platform-specific features, but rather on general interaction patterns between content producers and their audiences, making its application to other platforms–where visibility and influence may be structured through different mechanisms–a natural direction for future research.